Circular No. 9057

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2009 A8–A10 AND C/2009 O1 (STEREO) Further to IAUC 9054, additional Kreutz sungrazers have been found

on STEREO spacecraft images (CO = COR2-A and COR2-B). K. Battams, Naval Research Laboratory, notes that the three January comets were all small, faint, and slightly diffuse (mag around 11–12). C/2009 O1, which may have been around mag 6 or 6.5, appeared as a somewhat-long thin streak with no discernable head or coma in COR2-A images; in COR2-B images, the comet appeared much more condensed (possibly slightly diffuse and with a hint of a barely detectable tail).

Comet  2009 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2009 A8 Jan. 4.544 22^{h}19.^{m}2 −13^{o}58′ H1A AW 2009-O25
C/2009 A9  10.600 22 44.7 −11 41 H1A RM 2009-O25
C/2009 A10  10.794 22 43.2 −11 26 H1A AW 2009-O25
C/2009 O1 July 18.870 11 12.2 + 2 32 CO MK 2009-O25

COMET C/2009 O2 (CATALINA) An apparently asteroidal comet discovered in the course of the Catalina

Sky Survey (discovery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found by other CCD astrometrists to show cometary appearance. J. V. Scotti, Lunar and Planetary Laboratory, writes that images taken on July 30.2 UT with the Kitt Peak 1.8-m f/2.7 reflector shows the object to be slightly diffuse with an 8′′ coma and a faint tail 0′.35 long in p.a. 155^{o}. W. H. Ryan reports that his images from July 30.35–30.37 with the 2.4-m f/8.9 reflector at Magdalena Ridge Observatory show a coma and a very faint tail in p.a. ∼ 135^{o}.

2009 UT  α_{2000}  δ_{2000}  Mag.
July 27.29859  20^{h}44^{m}36.^{s}52  −11^{o}30′35′′.0  19.5

The available astrometry, the following preliminary parabolic orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2009-O53.

T = 2010 Mar. 24.656 TT  ω = 133.^{o}010
Ω = 310.266 }2000.0
q = 0.70282 AU  i = 108.284
2009 July 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green